OpenStudy (anonymous):

solve the equation for t, 0 less than or equal to t < 2pi 2cot(t)-csc(t)=3 choices are {1.18} {0.87,3.45} {0.87,3.14,3.45} {0.84,4.24}

OpenStudy (anonymous):

Simple cot(t)= cos(t)/sin(t) Csc(t)=1/sint 2cos(t)-1=3sin(t) Cos(t)=sqrt(1-(sin(t))^2)) 1-(sint)^2=((3sint+1)/2))^2 4-4(sint)^2=9(sint)^2+6sint+1 5(sint)^2+6sint-3=0 now you have a quadratic in terms of sine t solve it you would get two values for sin t now since the value of sint 0<=Sin t<=1 there for any value less than zero and any value greater than zero would be discarded

OpenStudy (anonymous):

now use the formula for quadratics to obtain the root becuase this equation cannot be factorised

OpenStudy (anonymous):

Above should be 13 sin ^2 t not 5 sin^2 t. Anyway, I think there is something wrong with the question...

OpenStudy (anonymous):

@estudier the question is correct look at my simplification

OpenStudy (anonymous):

Your simplification contains the above error. But either way, the answer choices or the question is wrong.

OpenStudy (anonymous):

@estudier right it would be 13

OpenStudy (anonymous):

after I put it in quadratic equation I get \[\frac{ -6\pm \sqrt{6^{2}-4(13)(-3)} }{ 2(13) }\] then I simplify to \[\frac{ -6\pm 8\sqrt{3} }{ 26 }\]

OpenStudy (anonymous):

that leaves me with \[\frac{ -3\pm 4\sqrt{3} }{ 13 }\]

OpenStudy (anonymous):

put 4sqrt(3)=4x1.732